### v2

# Significance of tensor force in pseudo-spin symmetry

Abstract
We point out that the pseudo-spin symmetry (PSS) of nuclei significantly depends on the proton ($Z$) and neutron numbers ($N$), sometimes giving rise to the characteristic structures. By using the non-relativistic spherical Hartree-Fock calculation with a realistic tensor force, we show that the tensor force may be deeply relevant to the $Z$- and $N$-dependence of the PSS. While the PSS has often been discussed in the context of the relativistic symmetry, the tensor-force effects on the PSS sometimes look analogous to the $Z$-and $N$-dependence of the PSS in the relativistic mean-field (RMF) calculations without explicit tensor force. The observed variation of the $p0d_{3/2}$-$p1s_{1/2}$ levels from 40Ca to 34Si is consistent with the tensor-force-driven $Z$-dependence of the PSS, but not necessarily with the RMF result. Even though it is too early to be conclusive, this result elucidates the significance of the tensor force when discussing the PSS.
Introduction.
Various symmetries have been found and discussed in nuclear structure physics. The pseudo-spin symmetry (PSS) ref:AHS69 ; ref:HA69 is among them, being crucial for some characteristic structures as discussed below. The so-called pseudo-LS coupling scheme and the pseudo-SU(3) model were also proposed on top of the PSS ref:AHS69 . In short, the PSS is the near degeneracy between $(n,\ell,j=\ell+1/2)$ and $(n^{\prime}=n-1,\ell^{\prime}=\ell+2,j^{\prime}=\ell^{\prime}-1/2)$ single-particle (s.p.) orbitals found in a certain region of the nuclear chart, where $(n,\ell,j)$ stands for the radial, orbital, and summed angular-momentum quantum numbers. For those pairs, the pseudo-radial and orbital quantum numbers are assigned as $\tilde{n}=n-1=n^{\prime}$ and ~=+1$\tilde{\ell}=\ell+1=\ell^{\prime}-1$. It has been argued since the late 1990s that the PSS is a relativistic symmetry ref:Gin97 ; ref:Men98 ; ref:LMZ15 , as $\tilde{\ell}$ is the real orbital angular momentum of the lower component of the Dirac spinor, rather than a hypothetical quantum number.
The tensor force has been pointed out to give rise to proton- ($Z$) and neutron-number ($N$) dependence of the shell structure ref:Vtn ; ref:Les07 ; ref:SC14 ; ref:NS14 , which is sometimes called “shell evolution”. This $Z$- and $N$-dependence of the shell structure should be relevant to the PSS. Whereas tensor-force effects on the PSS have been argued ref:WYLG13 , sufficient attention has not been paid to the $Z$- and $N$-dependence. In this paper, we discuss how the tensor force affects the PSS, with particular interest in the variation due to the occupation of specific orbits.
We apply the spherical Hartree-Fock (HF) calculations, and compare the s.p. energy spacings between the PSS partners among effective interactions (or energy-density functionals) with and without the tensor force. Because the tensor force is active in $jj$-closed configurations while inactive under the spin saturation ref:SNM16 , the variation of the spacings between a $\ell s$-closed nucleus and a $jj$-closed one provides a good indication of the tensor-force effects. In investigating the tensor-force effects, the M3Y-type semi-realistic interaction ref:Nak03 has a desirable property. It contains a realistic tensor force, which has been derived from a bare nucleonic interaction through the $G$-matrix ref:M3Y-P without further adjustment, and at the same time the strength of the tensor force is well examined in the shell structure ref:NSM13 . We calculate s.p. energies with the M3Y-P6 interaction ref:Nak13 ; ref:Nak20 , which provides the magic numbers compatible with most experimental data up to unstable nuclei ref:NS14 . They are compared with those of the Gogny-D1S interaction ref:D1S , which is a representative of the phenomenological interactions that have been used for a variety of nuclear structure studies with no explicit tensor force. To clarify the tensor-force effects, we also present the s.p. energies obtained with M3Y-P6 but after subtracting the contribution of the tensor force. The s.p. levels in the relativistic mean-field (RMF) approach with the DD-ME2 parameter-set ref:DD-ME2 are shown as well, for which the DIRHB code ref:DIRHB has been employed. Note that, under the time-reversal symmetry, the tensor force has contributions only from the exchange terms ref:SNM16 ; ref:VB72 ; ref:WZLL18 , which are not included in the RMF calculations. Experimental information on the s.p. levels can be obtained from the low-lying levels adjacent to the doubly-closed nuclei. Despite the fragmentation of the s.p. strengths, the energy difference between the lowest-lying states is often an acceptable measure, particularly within a single nuclide. We shall compare the calculated levels with the available experimental results.
$Z$- and $N$-dependence of pseudo-spin symmetry.
While the PSS plays important roles in nuclear structure, we here point out that the PSS has $Z$- and $N$-dependence, well maintained in certain regions but almost lost in others. In discussing the origin of the PSS, the $Z$- and $N$-dependence should not be discarded.
At the mean-field level, the tensor force has the following effects ref:Vtn ; ref:Les07 ; ref:SC14 ; ref:NS14 ; ref:SNM16 ; ref:Nak20 . It predominantly acts between a proton and a neutron. At the $\ell s$ closure, the tensor-force effects become negligibly small owing to the spin saturation. When a proton (neutron) orbit with $j=\ell+1/2$ is occupied, the tensor force pushes up neutron (proton) $j^{\prime}=\ell^{\prime}+1/2$ orbits and pushes down neutron (proton) $j^{\prime}=\ell^{\prime}-1/2$ orbits. Thus, the tensor force yields significant $Z$- and $N$-dependence of the s.p. level spacings through the occupation of specific orbitals.
We first look at the neutron shell structure above $N=28$. The near degeneracy of $n0f_{5/2}$ and $n1p_{3/2}$ in the Ni nuclei is a typical example of the PSS ref:AHS69 . On the other hand, the magicity of $N=32$ near 52Ca ref:Ca52_Ex2 occurs due to the large spacing between these orbits, implying that the PSS significantly depends on the proton number $Z$. In Fig. 1, the calculated s.p. energy spacing between the PSS partners at 48Ca is compared with the spacing at 56Ni. Occupation of protons on $p0f_{7/2}$ gives rise to the difference between 48Ca and 56Ni. This $Z$-dependence of the shell structure may largely be ascribed to the tensor force ref:Nak10b . The tensor force is active at 56Ni owing to the occupation on $p0f_{7/2}$, while its effect is small at 48Ca because of the $\ell s$-closure for protons. The $Z$-dependence in the M3Y-P6 result is compatible with the experimental levels extracted from the lowest levels at 49Ca and 57Ni. In contrast, D1S does not give the $Z$-dependence accounting for the PSS around the Ni nuclei. If we subtract the tensor-force contribution, M3Y-P6 reaches quite a similar result to that of D1S as shown by the dashed bars in Fig. 1, confirming that the tensor force is responsible for the $Z$-dependence. In all examples handled in this paper, the s.p. spacings with M3Y-P6 resemble those with D1S if subtracting the tensor-force contribution.
Figure 1: S.p. energy of $n0f_{5/2}$ (blue bars) measured from its PSS partner $n1p_{3/2}$ (orange bars) at 48Ca and 56Ni. The spherical HF results with D1S and M3Y-P6 are depicted. The result after subtracting the tensor-force contribution from that with M3Y-P6 is shown by the dashed bars, and the RMF result with DD-ME2 is also presented. Energies of $n0f_{7/2}$ (black bars) and $n1p_{1/2}$ (green bars) are also displayed for reference. For comparison, the measured energies of 49Ca and 57Ni of the lowest levels with corresponding spin-parities ref:NuDat are plotted. For $n0f_{7/2}$, the reference energy is extracted from the binding energies of 47Ca and 55Ni ref:AME2020 .
Intriguingly, the RMF result with DD-ME2 also describes the $Z$-dependence. Without the exchange terms, the origin of the $Z$-dependence in the DD-ME2 result should be attributed to the central channel. When we look at the calculated $n0f_{7/2}$ level together with $n0f_{5/2}$ and $n1p_{3/2}$, we find that both $n0f_{7/2}$ and $n0f_{5/2}$ shift down from 48Ca to 56Ni in the DD-ME2 result, almost keeping the spacing between them. Thus, the $Z$-dependence of the PSS in the DD-ME2 result originates from the different $Z$-dependence among the s.p. levels with different $\ell$ values, not greatly varying the $\ell s$ splitting. This is a sharp contrast to the M3Y-P6 result. Similar behavior is found in other regions shown below. If we extract the s.p. energies from the measured binding energies of 47Ca and 55Ni, the variation of the $n0f_{7/2}$ level is inbetween the M3Y-P6 and DD-ME2 results. Considering together the ambiguity in extracting the s.p.energies due to the fragmentation of the s.p. states, these data cannot conclusively tell the preference between the tensor-force picture and the relativistic picture of the $Z$-dependence.
The relative position of $n1p_{1/2}$ is also displayed in Fig. 1. The $1/2^{-}$ level lies closely to $3/2^{-}$ and $5/2^{-}$ experimentally. The position of the $1/2^{-}$ level does not look to be well reproduced at 56Ni in the D1S and M3Y-P6 results. However, our main concern is the $Z$-dependence of the level spacing. The spacing between $n1p_{3/2}$ and $n1p_{1/2}$ does not change much from 48Ca to 56Ni in all the results, although it slightly widens in the D1S and tensor-subtracted M3Y-P6 results, opposite to the experimental data.
Let us turn to other cases. Whereas the $p1s_{1/2}$ level lies well below its PSS partner $p0d_{3/2}$ near 40Ca, the two levels are inverted and closely lying around 48Ca, giving an example of the $N$-dependence of the PSS. It has been pointed out that the energy difference of these s.p. levels is governed by the tensor force ref:NSM13 . The role of the tensor force is again confirmed in Fig. 2, in which the variation of the energy spacing between $p0d_{3/2}$ and $p1s_{1/2}$ from 40Ca to 48Ca is shown. The tensor-force effect is sizable at 48Ca owing to the occupation on $n0f_{7/2}$, while small at 40Ca. The experimental levels are extracted from the lowest levels of 39K and 47K with corresponding spin-parity. The energy difference is similar to the one obtained from the proton knockout reaction data, which almost exhaust the spectroscopic factors ref:DWKM76 ; ref:Ogi87 . The variation of the difference between $p0d_{3/2}$ and $p1s_{1/2}$ is insufficient in the RMF result ref:GMK07 ; ref:WGZD11 . Notable difference is found between the full M3Y-P6 result and the others in the variation of $p0d_{5/2}$ from 40Ca to 48Ca, reflecting the tensor force, though the energy of this deep hole state is not easy to evaluate from experimental data because of its fragmentation.
Figure 2: S.p. energy of $p1s_{1/2}$ (orange bars) measured from its PSS partner $p0d_{3/2}$ (blue bars) at 40Ca and 48Ca. Energy of $p0d_{5/2}$ (black bars) is also displayed. The measured energies of 39K and 47ref:NuDat are shown for comparison. See Fig. 1 for other conventions.
The $E2$ excitation strengths in the neutron-deficient Sn nuclei are hardly described as far as the stiff spherical shape is assumed for their ground states ref:Tog18 ; ref:ONAT23 . The near degeneracy of the PSS partners $n0g_{7/2}$ and $n1d_{5/2}$ seems to trigger their deformability. In contrast, $N=56$ may be a submagic number at 96Zr ref:NS14 ; ref:MN18 , whose excitation energy is relatively high ref:NuDat . The $n0g_{7/2}$ level is distant from $n1d_{5/2}$ at Zr but becomes close as $p0g_{9/2}$ is occupied, providing another example of the $Z$-dependence of the PSS. We investigate the $Z$-dependence of the s.p. level spacing by picking up the $N=50$ nuclei. As observed in Fig. 3, M3Y-P6 well describes the experimental $Z$-dependence, owing to the tensor force. The level spacing between $n0g_{7/2}$ and $n1d_{5/2}$ seems too large at Sn with D1S and too small at Zr with DD-ME2. Although the variation of $n0g_{9/2}$ is distinguished between M3Y-P6 and the others, the data on 99Sn is required to access it experimentally, which lies outside the proton-drip line.
Figure 3: S.p. energy of $n0g_{7/2}$ (blue bars) measured from its PSS partner $n1d_{5/2}$ (orange bars) at 90Zr and 100Sn. Energy of $n0g_{9/2}$ (black bars) is also displayed. The measured energies of 91Zr and 101Sn ref:NuDat are shown for comparison. See Fig. 1 for other conventions.
Via the measurements of the isotope shifts, the differential charge radii have been extracted with high precision. Apart from exceptions like the neutron-deficient Sn nuclei mentioned above, the proton configuration hardly changes in magic-$Z$ isotopes, e.g., Pb, Sn, Ni and Ca. It has been disclosed that the charge radii prevalently have kinks at magic $N$ ref:Ang13 ; ref:shift-Ca52 ; ref:shift-Sn134 ; ref:Nak19 . For Pb, the near degeneracy of the PSS partners $n0i_{11/2}$ and $n1g_{9/2}$ was argued to be one of the key ingredients providing the kink at 208Pb ref:SLKR95 ; ref:RF95 ; ref:GSR13 ; ref:NI15 ; ref:shift-Hg . The near degeneracy of $n0h_{9/2}$ and $n1f_{7/2}$ could be important ref:Nak15 in the newly discovered kink at 132Sn ref:shift-Sn134 , as well. It is noted that not all RMF parameter-sets successfully reproduce the kink at 132Sn, possibly related to the degree of the PSS ref:NOSW23 . In Fig. 4, the variation of the energy spacing between $n0h_{9/2}$ and $n1f_{7/2}$ from 140Ce to 132Sn is presented. At 140Ce, $p0g_{7/2}$ is fully occupied in the calculated results, making the tensor-force effect on the neutron orbitals vanishingly small. As the observed excitation energy is relatively high ref:NuDat , $Z=58$ could be a submagic number at 140Ce ref:NS14 in reality. The $n0h_{9/2}$ orbit lies above $n1f_{7/2}$ at 140Ce, and it further goes up at 132Sn. However, the tensor force suppresses the increase of the $n0h_{9/2}$ energy, keeping the PSS a good approximation and producing a kink at 132Sn via a three-nucleon-force effect ref:Nak15 , whereas another study ascribes the origin of the kink to the property of the pairing ref:Fay00 . Analogously, the PSS maintained by the tensor force is important for the kink of the charge radii at 208Pb.
Figure 4: S.p. energy of $n0h_{9/2}$ (blue bars) measured from its PSS partner $n1f_{7/2}$ (orange bars) at 140Ce and 132Sn. Energies of $n0h_{11/2}$ (black bars) and $n2p_{3/2}$ (lavender bars) are also displayed. The measured energies of 141Ce and 133Sn ref:NuDat are shown for comparison. See Fig. 1 for other conventions.
Let us go back to the region near 40Ca. The $n1s_{1/2}$ orbit lies below its PSS partner $n0d_{3/2}$. It may go up and down as $p0d_{3/2}$ becomes unoccupied. D1S and DD-ME2, which do not contain explicit tensor force, predict $n1s_{1/2}$ goes down from 40Ca to 34Si, while M3Y-P6 gives the opposite trend. These qualitatively different predictions are noteworthy, because they seem distinctive of the tensor-force effect. The 34Si nucleus is expected to have a doubly-magic nature, with a good possibility of the proton semi-bubble density ref:NSM13 ; ref:LLSZ16 . Therefore, experimental data at 39Ca and 33Si could supply a reference. As shown in Fig. 5, the measured energies of the lowest levels are consistent with the M3Y-P6 result.
Figure 5: S.p. energy of $n1s_{1/2}$ (orange bars) measured from its PSS partner $n0d_{3/2}$ (blue bars) at 40Ca and 34Si. Energy of $n0d_{5/2}$ (black bars) is also displayed. The measured energies of 39Ca and 33Si ref:NuDat are shown for comparison. See Fig. 1 for other conventions.
We have discussed variations of the s.p. level spacings between the pseudo-spin partners through the occupation of specific orbits. One might expect that a long chain of isotopes or isotones gives clear indication of tensor-force effects in the PSS, possibly discriminating them from the relativistic effects. In Figs. 14, 15 and 16 of Ref. ref:Nak08b , the variation of the s.p. level spacing was investigated for the chains of $Z=50$, $N=50$ and $N=82$ nuclei, respectively. To elucidate the roles of the tensor force, results of the non-relativistic interactions with and without tensor force were compared. However, the central channels in the interaction, which are more or less adjusted to experimental data, also influence the level spacing, and obscure the tensor-force effects, whereas the tensor force significantly influences the relative energies of the unique-parity orbits. We have confirmed that this holds for the DD-ME2 case.
Summary and discussion.
We have pointed out the relevance of the $Z$- and $N$-dependence of the shell structure, often called “shell evolution”, to the pseudo-spin symmetry (PSS). This indicates that the tensor force may play a significant role in the PSS. We have investigated the tensor-force effects on the PSS by employing the spherical Hartree-Fock calculations with M3Y-P6. It should be noted that the tensor force is undoubtedly contained in the nucleonic interaction, and the tensor force adopted here is a realistic one derived from the $G$-matrix. We find that the tensor-force effects on the PSS sometimes look analogous to the $Z$-and $N$-dependence of the PSS in the relativistic mean-field (RMF) calculations without explicit tensor force, despite difference in the physics origin. However, for the variation of the $p0d_{3/2}$-$p1s_{1/2}$ levels from 40Ca to 34Si, the experimental data is compatible with the tensor-force-driven picture of the PSS, but not with the prediction of the RMF with the DD-ME2 Lagrangian. Although it is too early to conclude from this case that the tensor force dominates the $Z$- and $N$-dependence of the PSS, it demonstrates that the tensor force may be significant in the PSS. Conversely, it could be misleading to discuss the PSS only within the context of relativistic symmetry without paying attention to the tensor-force effects.
The tensor force between nucleons arises in the relativistic Hartree-Fock (RHF) framework ref:BMGM87 ; ref:Long07 , which takes account of the exchange terms. The role of the tensor coupling due to the $\rho$-meson in the PSS has been argued ref:Geng19 , which gives a part of the tensor-force effect discussed here. See Ref. ref:WZLL18 for the relation of the meson-nucleon coupling in the RHF to the nucleonic tensor force. As the tensor force and the relativistic effects may affect the PSS cooperatively and competitively, it is of interest to view how the $Z$- and $N$-dependence of the PSS arises within the RHF scheme.
Discussions with N. Hinohara and T. Naito are gratefully acknowledged. This work is supported by the JSPS KAKENHI, Grant No. JP24K07012. A part of the numerical calculations has been performed on HITAC SR24000 at Institute of Management and Information Technologies at Chiba University.